izukumidoriya:

what is 4/5.6 divided bye 43/4.7?

Downpour17:

do you have a calculator?

blakev1503:

it is 0.0780730897

izukumidoriya:

oh thank you and i do not have a calculator

blakev1503:

your welcome

CripQUEZZ:

When u divide fractions u use kcf u keep the first fraction change the division to multiplication sign and flip the second fraction around to 4.7/43 and then u multiply across and simplify and then u get ur answer

jhonyy9:

so use this \[\left( \frac{ \frac{ a }{ b }}{ \frac{ c }{ d } } \right) = \left( \frac{ a }{ b } \right)*\left( \frac{ d }{ c } \right)\]

CripQUEZZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @blakev1503 it is 0.0780730897 \(\color{#0cbb34}{\text{End of Quote}}\) It’s supposed to be a fraction not a decimal and u cant give direct answers

CripQUEZZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 so use this \[\left( \frac{ \frac{ a }{ b }}{ \frac{ c }{ d } } \right) = \left( \frac{ a }{ b } \right)*\left( \frac{ d }{ c } \right)\] \(\color{#0cbb34}{\text{End of Quote}}\) Yes thank u sum1 who understands

izukumidoriya:

thank you all

CripQUEZZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @izukumidoriya thank you all \(\color{#0cbb34}{\text{End of Quote}}\) Np !